Question 1:

Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,600 helmets, using 2,304 kilograms of plastic. The plastic cost the company $17,510.

According to the standard cost card, each helmet should require 0.55 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,600 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,600 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?

Question 2:

SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 4,300 of these meals using 1,250 direct labor-hours. The company paid its direct labor workers a total of $11,250 for this work, or $9.00 per hour.

According to the standard cost card for this meal, it should require 0.30 direct labor-hours at a cost of $8.50 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 4,300 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 4,300 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

Question 3:

Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.50q
Indirect labor $4,000 + $1.40q
Utilities $5,000 + $0.60q
Supplies $1,200 + $0.30q
Equipment depreciation $18,100 + $2.90q
Factory rent $8,600
Property taxes $2,800
Factory administration $13,100 + $0.70q

The Production Department planned to work 4,100 labor-hours in March; however, it actually worked 3,900 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 65,970
Indirect labor $ 8,920
Utilities $ 7,850
Supplies $ 2,640
Equipment depreciation $ 29,410
Factory rent $ 9,000
Property taxes $ 2,800
Factory administration $ 15,200

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Calculate the spending variances for all expense items.

Question 4:

Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below:

Standard Quantity or Hours Standard Price
or Rate
Standard
Cost
Direct materials 7.90 pounds $ 2.10 per pound $ 16.59
Direct labor 0.50 hours $ 5.00 per hour $ 2.50

During the most recent month, the following activity was recorded:

  1. Fourteen thousand eight hundred and fifity pounds of material were purchased at a cost of $2.00 per pound.

  2. All of the material purchased was used to produce 1,500 units of Zoom.

  3. 600 hours of direct labor time were recorded at a total labor cost of $4,200.

Required:

1. Compute the materials price and quantity variances for the month.

2. Compute the labor rate and efficiency variances for the month.

Question 5:

Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below:

Standard Quantity or Hours Standard Price
or Rate
Standard
Cost
Direct materials 7.00 pounds $ 3.00 per pound $ 21.00
Direct labor 0.40 hours $ 14.00 per hour $ 5.60


During the most recent month, the following activity was recorded:

  1. Thirteen thousand six hundred pounds of material were purchased at a cost of $2.90 per pound.

  2. The company produced only 1,360 units, using 12,240 pounds of material. (The rest of the material purchased remained in raw materials inventory.)

  3. 644 hours of direct labor time were recorded at a total labor cost of $7,728.

Required:

Compute the materials price and quantity variances for the month. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

1.

Standard Quantity allowed = 3600*0.55

= 1980

Standard cost allowed = 1980*8

= 15840

Materials spending variance = 15840 - 17510

= 1670 U

Materials Price Variance = (SP-AP) *AQ = (8 - 17510/2304)*2304

= 922 F

Materials Quantity Variance = (SQ-AQ) *SP = (1980-2304)*8

= 2592 U